OpenStudy (anonymous):

A force F acts in the x - direction, its magnitude given by F = ax^2, where x is in meters and a =5.0N/m^2 . Find the work done by this force as it acts on a particle moving from x = 0 to x = 6.0 m.

OpenStudy (anonymous):

F = ax^2 plug in 0, and 6 which gives you 180J and 0J, subtract them and you get 180J

OpenStudy (anonymous):

W=F(d) = 180(6) = 1080 watts

OpenStudy (anonymous):

lol for reals?

OpenStudy (anonymous):

Pretty sure broski

OpenStudy (anonymous):

did you finish the whole thing?

OpenStudy (anonymous):

Currently on #4

OpenStudy (anonymous):

Wow! Your gettin it done early!

OpenStudy (anonymous):

;) go on skype and we can talk